OpenStudy (anonymous):

Solve the following system of equations. 2x + y = 3 x = 2y - 1

OpenStudy (anonymous):

plug in x=2y-1 to the first equation.. 2(2y-1)+y=3 4y-2+y=3 5y-2=3 5y-2+2=3+2 5y=5 y=1 and then plug in y to an equation and itll give you x

HanAkoSolo (jamierox4ev3r):

@Zacharymartin do you still need help on this question?

OpenStudy (anonymous):

Oh no i forgot to close it. My bad

HanAkoSolo (jamierox4ev3r):

haha okay awesome :D